# Conveyor-belt superconducting quantum computer

### v1

@@ Abstract
The processing unit of a solid-state quantum computer consists in an array of coupled qubits, each locally driven with on-chip microwave lines that route carefully-engineered control signals to the qubits in order to perform logical operations. This approach to quantum computing comes with two major problems. On the one hand, it greatly hampers scalability towards fault-tolerant quantum computers, which are estimated to need a number of qubits – and, therefore driving lines – on the order of $10^{6}$. On the other hand, these lines are a source of electromagnetic noise, exacerbating frequency crowding and crosstalk, while also contributing to power dissipation inside the dilution fridge. We here tackle these two overwhelming challenges by presenting a novel quantum processing unit (QPU) for a universal quantum computer which is globally (rather than locally) driven. Our QPU relies on a string of superconducting qubits with always-on ZZ interactions, enclosed into a closed geometry, which we dub “conveyor belt”. Strikingly, this architecture requires only $\mathcal{O}(N)$ physical qubits to run a computation on $N$ computational qubits, in contrast to previous $\mathcal{O}(N^{2})$ proposals for global quantum computation. Additionally, universality is achieved via the implementation of single-qubit gates and a one-shot Toffoli gate. The ability to perform multi-qubit operations in a single step could vastly improve the fidelity and execution time of many algorithms.
@@ I Introduction
Scalability in Quantum Computing (QC) refers to the ability to increase the number of qubits and other resources in a quantum system while maintaining performance, error rates, and coherence times Fowler2012 . Achieving scalability is essential for building quantum computers capable of solving complex problems beyond the reach of classical computers. For devices based on superconducting circuits Gambetta2017 ; Martinis2019quantum ; SQ2020 ; supremacy_PRL_2021 ; blais_review_RMP ; bravyi2022future ; ezratty2023perspective (and other solid-state platforms), scalability is fundamentally constrained, among others, by the need for localized drive of each physical qubit You2005supercond ; Girvin2008 ; devoret2013supercond ; wedin2017quantum in order to perform single-qubit rotations and entangling operations as required by the computation. This issue, often called the “wiring problem”, leads to a wiring overload Gambetta2017 ; tsai2020pseudo ; tsai2021gate : for example, if we had to scale the current microwave interface to control a million-qubit system, the electronics itself would take up to three football fields of space, while consuming about 40 MW of dc power Bardin2020 . A potential solution to this problem is the development of globally driven QC schemes Lloyd_1993 ; benjamin_2000 ; benjamin_2001 ; benjamin_2003 ; benjamin_2004 . These schemes involve designing and operating quantum computers in a manner that minimizes or eliminates the need for precise, localized manipulation of individual qubits. The main benefits include the simplification of designs due to the reduced complexity of control electronics and interconnections, and potentially enhanced error resilience with respect to local errors. However, globally driven QC approaches also face challenges, such as ensuring precision in global operations, creating and maintaining large, possibly highly entangled states for logical information encoding, and implementing fault-tolerant operations globally, which require sophisticated error-correcting protocols.
@@
A recently proposed globally driven superconducting QC architecture menta2024globally addresses some of these challenges, specifically the precision in global operations and effectiveness of state preparation. Building upon a proposal based on Rydberg atoms cesa2023universal ; fromonteil2024hamilton , in Ref. menta2024globally we demonstrated that a globally driven universal quantum computer can be achieved in a superconducting platform by leveraging the longitudinal (always-on) ZZ interaction ZZ-suppression2020 ; Xu2024 ; ZZ-contrast2020 ; ZZ-eliminating2022 ; ZZ_2020 ; ZZ_2021 ; xu2021z ; Fors2024 between neighbouring superconducting qubits organized in a two-dimensional (2D) ladder. In this work we present a novel architecture that significantly improves the design proposed in Ref. menta2024globally by drastically reducing the resources needed to implement it.
@@
Figure 1: Schematic description of the proposed architecture. Two types, $A$ and $B$, of superconducting qubits (red and blue squares, respectively) are separately driven by two classical sources $V_{A,B}(t)$ (red and blue continuous lines). They are coupled via a longitudinal ZZ coupling (black and grey springs). Black triangles inside regular and crossed qubits denote local corrections in the resonance frequency of the qubits (see main text). The $A$-type crossed qubit (red square inside the loop) enables one-shot Toffoli gate (three-qubit gate) – the corresponding interactions are depicted in gray. The $B$-type crossed qubit performs single-qubit gates. The elements highlighted in yellow indicate the information carrying sites $Q_{1}$, $Q_{2}$, $\cdots$, $Q_{N}$, separated by three-qubits sectors $S_{1}$, $S_{2}$, $\cdots$, $S_{N}$. The $Q_{j}$’s host the computational qubits through well-formed configurations (2) where the $S_{j}$’s are initialized in alternating sequences of paramagnetic ($geg$) and ferromagnetic ($ggg$) phases. Since ZZ interactions critically depend on the relative distance between two qubits, we have used the honeycomb structure to make sure that the couplings among adjacent qubits are (nominally) all identical. The figure pertains to a $N=8$ qubit quantum computer.
@@
At variance with previous schemes menta2024globally ; cesa2023universal , which require to operate globally on a 2D array of ${\cal O}(N^{2})$ physical qubits to run universal operations on $N$ computational qubits, the setup discussed here only needs ${\cal O}(N)$ elements – specifically $14N+$ 1 – organized in a closed loop, as shown in Fig. 1. For clarity, we stress that the computational qubits are the ones in which we encode the logical information. On the other hand, the set of physical qubits encompasses all two-level systems present in the architecture. Sequences of control pulses that act collectively on half of the elements of the device allow us to induce special permutations on the states of the sites. Combining these transformations with the possibility to implement both single-qubit gates and a one-shot Toffoli gate yields universal quantum computation. It is important to mention that the associated dynamical processes are based on a blockade regime, analogous to the one in Rydberg atoms Rydberg_blockade , which is here achieved thanks to a strong ZZ coupling between nearest-neighbor qubits. The main features of this effect for the case of superconducting qubits are described in Ref. menta2024globally . We also mention that, in QC, the ability to perform multi-qubit operations, such as the Toffoli gate, in a single step could vastly improve the fidelity and execution time of many algorithms Zahedinejad2015 ; baker2022 . To the best of our knowledge, our proposal introduces a novel superconducting platform that, for the first time, incorporates a one-shot Toffoli gate Nielsen2010 , which operates based on the ZZ-induced blockade regime.
II Results
II.1 Model
@@
The system shown in Fig. 1 consists of a closed loop formed by $4N$ superconducting qubits coupled via a spatially-uniform ZZ interaction of coupling strength $\hbar\zeta$ (black springs in the figure). The loop contains $N$ sites, identified by the symbols $Q_{1}$, $Q_{2}$, $\cdots$, $Q_{N}$, which play a special role in the model and will be referred to as Information Carrying (IC) sites. For all $j\in\{1,\cdots,N\}$ the IC site $Q_{j}$ is separated from its adjacent counterpart $Q_{j+1}$ by a sector $S_{j}$ which contains three ”non-active” sites. In the figure, there is an extra superconducting qubit placed inside the loop, which is connected to the first three IC sites $Q_{1}$, $Q_{2}$, and $Q_{3}$ via the same ZZ coupling term that connects every other element (these three couplings are highlighted in a light gray color to emphasize the fact that this additional qubit is placed inside the loop and not along it). The superconducting qubits in the model belong to two distinct families: $A$-type qubits, represented as red squares, and $B$-type qubits, represented as blue squares, which alternate in the loop in an $ABA$ pattern. As shown in the figure, all the IC sites host a $B$-type qubit, while the intermediate sectors $S_{j}$ are formed by one $B$-type and two $A$-type qubits. All qubits of the same type share the same level spacing $\hbar\omega_{A/B}$, except for the elements marked with triangles, which, for reasons that will be clear in the following, require $\zeta$-dependent local corrections to their level spacings i.e. $\hbar\omega_{A/B}\to\hbar(\omega_{A/B}+\zeta)$, due to the higher number of nearest neighbors they have (three) compared to the other qubits in the setup. To eliminate the possibility of a residual “swap” term between nearest-neighbor qubits, it may be necessary to have $\omega_{A}\neq\omega_{B}$. Each family of qubits is collectively driven by the same time-dependent external source, termed $V_{A/B}(t)$. Thus, the control is global. The device also includes two inhomogeneities: the additional $A$-type element previously mentioned and one of the three $B$-type elements connected to it (specifically, the one located at the IC site $Q_{2}$). These special qubits, referred to as crossed qubits, are used to perform multi- and single-qubit operations. They maintain the nominal level spacing of their respective families but have an augmented coupling with the external source, i.e. the Rabi frequency. As shown in Ref. menta2024globally , this difference allows independent control of the normal qubits and crossed qubits of a given family, despite the global control pulse acting simultaneously on them. Both the $A$- and $B$-type crossed qubits are marked with a small black triangle. Finally, a third control line, $V_{\mathrm{init}}(t)$, acting on a subset of the $B$-type qubits, is necessary for the initialization of the machine. This control line is active only at the initial stage of the computation. Regarding the readout procedure, a possible implementation is discussed in the supplemental material supplemental .
@@
Adopting the same convention of Ref. menta2024globally we write the Hamiltonian of the setup as $\hat{H}(t):=\hat{H}_{0}+\hat{H}_{\rm drive}(t)$, where $\hat{H}_{0}:=\sum_{\chi\in\{A,B\}}\sum_{i\in\chi}\frac{\hbar\omega_{i}}{2}\hat{\sigma}_{i}^{(z)}+\sum_{\langle i,j\rangle}\frac{\hbar\zeta}{2}\hat{\sigma}_{i}^{(z)}\otimes\hat{\sigma}_{j}^{(z)}$, describes the local energy contribution of the superconducting qubits and their ZZ interactions which are fixed by the geometry of the model, while $\hat{H}_{\rm drive}(t):=\sum_{\chi\in\{A,B\}}\sum_{i\in\chi}\hbar\Omega_{\chi}(t)\sin(\omega_{\mathrm{d},\chi}t+\phi_{\chi}(t))\hat{\sigma}_{i}^{(y)}$, is the time-dependent driving contribution induced by the classical control lines. In these equations, $\hat{\sigma}_{i}^{(x,y,z)}$ represent ordinary Pauli matrices acting on the Hilbert space of the $i$-th qubit, expressed in the local energy basis $\{|g_{i}\rangle,|e_{i}\rangle\}$. The summation in the interacting part of $\hat{H}_{0}$ encompasses all nearest-neighbor interactions, mathematically representing all the black springs in Fig. 1. The parameter $\omega_{\mathrm{d},\chi}$ denotes the oscillation frequency of the driving pulse $V_{\chi}(t)$, while $\Omega_{\chi}(t)$ and $\phi_{\chi}(t)$ define the time-dependent Rabi frequency and phase of such control. In the following, we will assume that these quantities assume constant values on disjoint time windows, in such a way that, at each time, only one species (either $A$ or $B$) is driven. For simplicity, in writing $\hat{H}_{\rm drive}(t)$, we have omitted the driving term associated with the control line $V_{\rm init}(t)$, which operates solely at the very beginning of the computational process on a specific subset of the physical qubits. Additionally, we did not explicitly highlight that when the site index $i$ in the expression for $\hat{H}_{0}$ identifies a $B$-type qubit marked with a black triangle, the corresponding qubit’s level spacing becomes $\hbar(\omega_{B}+\zeta)$ instead of $\hbar\omega_{B}$. The same substitution must be performed when the site index $i$ identifies an $A$-type qubit marked with a black triangle, i.e. an $A$-type crossed qubit. Similarly, whenever in the expression of $\hat{H}_{\rm drive}(t)$ the index $i$ identifies a crossed qubit, $\Omega_{\chi}(t)$ needs to be replaced by $2\Omega_{\chi}(t)$. Apart from these local adjustments (which need to be engineered once and for all at fabrication level), it is important to note that $\Omega_{\chi}(t)$, $\phi_{\chi}(t)$, and $\omega_{\mathrm{d},\chi}$ are independent of the site index $i$, indicating that they are associated with a control pulse acting globally on all qubits of $\chi$-type in the model. We conclude this Section by stressing that the above Hamiltonian is exactly the same as the one studied in Ref. menta2024globally . However, due to the more exotic geometry, the dynamical features of the system are more complex than in the case of the 2D ladder geometry used in Ref. menta2024globally . We take advantage of this complexity to achieve the $\mathcal{O}(N)$ scaling, as opposed to the $\mathcal{O}(N^{2})$ scaling of Refs. menta2024globally ; cesa2023universal . In the next paragraph we describe such features. Further details can be found in the supplemental material supplemental as well as in Ref. menta2024globally .
II.2 Dynamical features
The first ingredient of the present proposal is to emulate the Rydberg blockade effect Rydberg_blockade using the ZZ interactions of the model. Specifically, one can show that under the strong-coupling condition $\eta_{\rm BR}:=|\zeta/\Omega_{\chi}|\gg 1$, by properly detuning the driving frequencies $\omega_{\mathrm{d},\chi}$ from the nominal level spacing of the qubits, one is able to selectively induce transitions among states of the system only when two nearest-neighbour sites do not simultaneously occupy their excited levels. The second fundamental ingredient is the presence of two different values of the Rabi frequency for qubits of the same family, i.e. the presence of regular and crossed qubits. Crucially, one can prove that this difference is sufficient to drive independently these two types of qubits even though their control is global menta2024globally . Specifically, one is able to perform generic unitaries defined as
$\hat{W}_{\chi}(\theta^{\prime},\bm{n}^{\prime};\theta^{\prime\prime},\bm{n}^{\prime\prime}):=\hat{W}_{\chi^{\rm r}}(\theta^{\prime},\bm{n}^{\prime})\hat{W}_{\chi^{\times}}(\theta^{\prime\prime},\bm{n}^{\prime\prime})~{},$ (1)

### v2

@@ Abstract
The processing unit of a solid-state quantum computer consists in an array of coupled qubits, each locally driven with on-chip microwave lines that route control signals to the qubits in order to perform logical operations. This approach to quantum computing comes with two major problems. On the one hand, it greatly hampers scalability towards fault-tolerant quantum computers, which are estimated to need a number of qubits – and, therefore driving lines – on the order of $10^{6}$. On the other hand, these lines are a source of electromagnetic noise, exacerbating frequency crowding and crosstalk, while also contributing to power dissipation inside the dilution fridge. We here tackle these two challenges by presenting a novel quantum processing unit (QPU) for a universal quantum computer which is globally (rather than locally) driven. Our QPU relies on a closed loop of superconducting qubits with always-on ZZ interactions which we dub “conveyor belt”. Notably, this architecture requires only $\mathcal{O}(N)$ physical qubits to run a computation on $N$ computational qubits, in contrast to previous $\mathcal{O}(N^{2})$ proposals for global quantum computation. Universality is achieved via the implementation of single-qubit gates and a one-shot Toffoli gate. The ability to perform multi-qubit operations in a single step could vastly improve the fidelity and execution time of many algorithms.
@@ I Introduction
Scalability in Quantum Computing (QC) refers to the ability to increase the number of qubits and other resources in a quantum system while maintaining performance, error rates, and coherence times Mohseni2024 ; google2025 ; Fowler2012 . Achieving scalability is essential for building quantum computers capable of solving complex problems beyond the reach of classical computers. For devices based on superconducting circuits Gambetta2017 ; Martinis2019quantum ; SQ2020 ; supremacy_PRL_2021 ; blais_review_RMP ; bravyi2022future ; ezratty2023perspective (and other solid-state platforms), scalability is fundamentally constrained, among others, by the need for localized drive of each physical qubit You2005supercond ; Girvin2008 ; devoret2013supercond ; wedin2017quantum in order to perform single-qubit rotations and entangling operations as required by the computation. This issue, often called the “wiring problem”, leads to a wiring overload Gambetta2017 ; tsai2020pseudo ; tsai2021gate : for example, if we had to scale the current microwave interface to control a million-qubit system, the electronics itself would take up to three football fields of space, while consuming about 40 MW of dc power Bardin2020 . A potential solution to this problem is the development of globally driven QC schemes Lloyd_1993 ; benjamin_2000 ; benjamin_2001 ; benjamin_2003 ; benjamin_2004 . These schemes involve designing and operating quantum computers in a manner that minimizes or eliminates the need for precise, localized manipulation of individual qubits. The main benefits include the simplification of designs due to the reduced complexity of control electronics and interconnections, and potentially enhanced error resilience with respect to local errors. However, globally driven QC approaches also face challenges, such as ensuring precision in global operations, creating and maintaining large, possibly highly entangled states for logical information encoding, and implementing fault-tolerant operations globally, which require sophisticated error-correcting protocols.
@@
A recently proposed globally driven superconducting QC architecture menta2024globally addresses some of these challenges, specifically the precision in global operations and the effectiveness of state preparation. Building upon a proposal based on Rydberg atoms cesa2023universal ; fromonteil2024hamilton , in Ref. menta2024globally , we demonstrated that a globally driven universal quantum computer can be achieved on a superconducting platform by leveraging the longitudinal (always-on) ZZ interaction ZZ-suppression2020 ; Xu2024 ; ZZ-contrast2020 ; ZZ-eliminating2022 ; ZZ_2020 ; ZZ_2021 ; xu2021z ; Fors2024 between neighboring superconducting qubits organized in a two-dimensional (2D) ladder. In this work, we present a novel architecture that significantly improves the design proposed in Ref. menta2024globally by drastically reducing the resources needed to implement it.
@@
Figure 1: Schematic description of the proposed architecture. Two types, AA and $B$, of superconducting qubits (red and blue squares, respectively) are separately driven by two classical sources $V_{A,B}(t)$ (red and blue continuous lines). They are coupled via a longitudinal ZZ coupling (black and grey springs). Black triangles inside regular and crossed qubits denote local corrections in the resonance frequency of the qubits (see main text). The $A$-type crossed qubit (red square inside the loop) enables one-shot Toffoli gate (three-qubit gate) – the corresponding interactions are depicted in gray. The $B$-type crossed qubit performs single-qubit gates. The elements highlighted in yellow indicate the information carrying sites $Q_{1}$, $Q_{2}$, $\cdots$, $Q_{N}$, separated by three-qubits sectors $S_{1}$, $S_{2}$, $\cdots$, $S_{N}$. The $Q_{j}$’s host the computational qubits through well-formed configurations (8) where the $S_{j}$’s are initialized in alternating sequences of paramagnetic (geggeg) and ferromagnetic (gggggg) phases. Since ZZ interactions critically depend on the relative distance between two qubits, we have used the honeycomb structure to make sure that the couplings among adjacent qubits are (nominally) all identical. The figure pertains to a $N=8$ qubit quantum computer.
@@
At variance with previous schemes menta2024globally ; cesa2023universal , which require to operate globally on a 2D array of ${\cal O}(N^{2})$ physical qubits to run universal operations on $N$ computational qubits, the setup discussed here only needs ${\cal O}(N)$ elements – specifically $4N+1$ – organized in a closed loop, as shown in Fig. 1. For clarity, we stress that the computational qubits are the ones in which we encode the logical information. On the other hand, the set of physical qubits encompasses all two-level systems present in the architecture. Sequences of control pulses that act collectively on half of the elements of the device allow us to induce special permutations on the states of the sites. Combining these transformations with the possibility to implement both single-qubit gates and a one-shot Toffoli gate yields universal quantum computation. It is important to mention that the associated dynamical processes are based on a blockade regime, analogous to the one in Rydberg atoms Rydberg_blockade , which is here achieved thanks to a strong ZZ coupling between nearest-neighbor qubits. The main features of this effect for the case of superconducting qubits are described in Ref. menta2024globally . We also mention that, in QC, the ability to perform multi-qubit operations, such as the Toffoli gate, in a single step could vastly improve the fidelity and execution time of many algorithms Zahedinejad2015 ; baker2022 . To the best of our knowledge, our proposal introduces a novel superconducting platform that, for the first time, incorporates a one-shot Toffoli gate Nielsen2010 , which operates based on the ZZ-induced blockade regime.
II Model
@@
The system shown in Fig. 1 consists of a closed loop formed by $4N$ superconducting qubits coupled via a spatially-uniform ZZ interaction of coupling strength $\hbar\zeta$ (black springs in the figure). The loop contains $N$ sites, identified by the symbols $Q_{1}$, $Q_{2}$, $\cdots$, $Q_{N}$, which play a special role in the model and will be referred to as Information Carrying (IC) sites. For all $j\in\{1,\cdots,N\}$ the IC site $Q_{j}$ is separated from its adjacent counterpart $Q_{j+1}$ by a sector $S_{j}$ which contains three “non-active” sites. In the figure, there is an extra superconducting qubit placed inside the loop, which is connected to the first three IC sites $Q_{1}$, $Q_{2}$, and $Q_{3}$ via the same ZZ coupling term that connects every other element (these three couplings are highlighted in a light gray color to emphasize the fact that this additional qubit is placed inside the loop and not along it). The superconducting qubits in the model belong to two distinct families: $A$-type qubits, represented as red squares, and $B$-type qubits, represented as blue squares, which alternate in the loop in an $ABA$ pattern. As shown in the figure, all the IC sites host a $B$-type qubit, while the intermediate sectors $S_{j}$ are formed by one $B$-type and two $A$-type qubits. All qubits of the same type share the same level spacing $\hbar\omega_{A/B}$, except for the elements marked with triangles, which, for reasons that will be clear in the following, require $\zeta$-dependent local corrections to their level spacings i.e. $\hbar\omega_{A/B}\to\hbar(\omega_{A/B}+\zeta)$, due to the higher number of nearest neighbors they have (three) compared to the other qubits in the setup. To eliminate the possibility of a residual “swap” term between nearest-neighbor qubits, it may be necessary to have $\omega_{A}\neq\omega_{B}$. Each family of qubits is collectively driven by the same time-dependent external source, termed $V_{A/B}(t)$. Thus, the control is global. The device also includes two inhomogeneities: the additional $A$-type element previously mentioned and one of the three $B$-type elements connected to it (specifically, the one located at the IC site $Q_{2}$). These special qubits, referred to as crossed qubits, are used to perform multi- and single-qubit operations. They maintain the nominal level spacing of their respective families but have an augmented coupling with the external source, i.e. the Rabi frequency. As proved in Ref. menta2024globally and summarized in Section III and more in detail in Appendix A, this difference allows independent control of the normal qubits and crossed qubits of a given family, despite the global control pulse acting simultaneously on them. Both the $A$- and $B$-type crossed qubits are marked with a small black triangle. Finally, a third control line, $V_{\mathrm{init}}(t)$, acting on a subset of the $B$-type qubits, is necessary for the initialization of the machine. This control line is active only at the initial stage of the computation. Regarding the readout procedure, a possible implementation is discussed in Appendix D.
Adopting the same convention of Ref. menta2024globally we write the Hamiltonian of the setup as $\hat{H}(t):=\hat{H}_{0}+\hat{H}_{\rm drive}(t)$, where
$\hat{H}_{0}:=\sum_{\chi\in\{A,B\}}\sum_{i\in\chi}\frac{\hbar\omega_{i}}{2}\hat{\sigma}_{i}^{(z)}+\sum_{\langle i,j\rangle}\frac{\hbar\zeta}{2}\hat{\sigma}_{i}^{(z)}\otimes\hat{\sigma}_{j}^{(z)}\;,$ (1)
describes the local energy contribution of the superconducting qubits and their ZZ interactions which are fixed by the geometry of the model, while
$\hat{H}_{\rm drive}(t):=\sum_{\chi\in\{A,B\}}\sum_{i\in\chi}\hbar\Omega_{\chi}(t)\sin(\omega_{\mathrm{d},\chi}t+\phi_{\chi}(t))\hat{\sigma}_{i}^{(y)}\;,$ (2)
@@
is the time-dependent driving contribution induced by the classical control lines. In these equations, $\hat{\sigma}_{i}^{(x,y,z)}$ represent ordinary Pauli matrices acting on the Hilbert space of the $i$-th qubit, expressed in the local energy basis $\{|g_{i}\rangle,|e_{i}\rangle\}$. The summation in the interacting part of $\hat{H}_{0}$ encompasses all nearest-neighbor interactions, mathematically representing all the black springs in Fig. 1. The parameter $\omega_{\mathrm{d},\chi}$ denotes the oscillation frequency of the driving pulse $V_{\chi}(t)$, while $\Omega_{\chi}(t)$ and $\phi_{\chi}(t)$ define the time-dependent Rabi frequency and phase of such control. In the following, we will assume that these quantities assume constant values on disjoint time windows, in such a way that, at each time, only one species (either $A$ or $B$) is driven. For simplicity, in writing $\hat{H}_{\rm drive}(t)$, we have omitted the driving term associated with the control line $V_{\rm init}(t)$, which operates solely at the very beginning of the computational process on a specific subset of the physical qubits. Additionally, we did not explicitly highlight that when the site index $i$ in the expression for $\hat{H}_{0}$ identifies a $B$-type qubit marked with a black triangle, the corresponding qubit’s level spacing becomes $\hbar(\omega_{B}+\zeta)$ instead of $\hbar\omega_{B}$. The same substitution must be performed when the site index $i$ identifies an $A$-type qubit marked with a black triangle, i.e. an $A$-type crossed qubit. Similarly, whenever in the expression of $\hat{H}_{\rm drive}(t)$ the index $i$ identifies a crossed qubit, $\Omega_{\chi}(t)$ needs to be replaced by $2\Omega_{\chi}(t)$. Apart from these local adjustments (which need to be engineered once and for all at fabrication level), it is important to note that $\Omega_{\chi}(t)$, $\phi_{\chi}(t)$, and $\omega_{\mathrm{d},\chi}$ are independent of the site index $i$, indicating that they are associated with a control pulse acting globally on all qubits of $\chi$-type in the model. We conclude this paragraph by stressing that the above Hamiltonian is exactly the same as the one studied in Ref. menta2024globally . However, due to the more exotic geometry, the dynamical features of the system are more complex than in the case of the 2D ladder geometry used in Ref. menta2024globally . We take advantage of this complexity to achieve the $\mathcal{O}(N)$ scaling, as opposed to the $\mathcal{O}(N^{2})$ scaling of Refs. menta2024globally ; cesa2023universal .
III Dynamical features and global control
The present proposal relies on two key ingredients. The first is the emulation of a Rydberg-blockade-like mechanism Rydberg_blockade through the $ZZ$ interactions between qubits. The second is the presence of two distinct Rabi frequencies within each qubit family, corresponding to regular and crossed qubits, which enables independent control despite being driven by the same global control.
III.1 Effective blockade dynamics
We consider global pulses $(\Omega_{\chi}(t),\phi_{\chi}(t))$ with $\chi=\{A,B\}$, applied piecewise in time such that only one qubit species is driven at any given instant. One can show that under the strong-coupling condition $\eta_{\rm BR}:=|\zeta/\Omega_{\chi}|\gg 1$, by properly detuning the driving frequencies $\omega_{\mathrm{d},\chi}$ from the nominal level spacing of the qubits, one is able to selectively induce transitions among states of the system only when two nearest-neighbour sites do not simultaneously occupy their excited levels. As a result, the dynamics is effectively restricted to configurations in which no two interacting neighbors occupy the excited state, realizing a blockade mechanism analogous to that of Rydberg systems.
In this limit, the unitary evolution $\hat{U}_{\chi}$ generated by a single global pulse of length $\tau$ acting on qubits of type $\chi$ can be written as menta2024globally